# Geometric Bounds for Persistence

### v1

@@ 2 Background
In this section, we overview the basic notions from applied algebraic topology and metric geometry used in the remainder of the paper. We give a summary of our notations in Table 1 in the Appendix.
@@ 2.1 Persistent Homology
Persistent Homology (PH) is a methodology with roots in Applied Algebraic Topology which helps capturing different characteristics of a given metric space. In the past two decades, PH has been employed as a powerful mathematical machinery for discovering patterns in data in applications within Machine Learning and Data Science. For more details on PH and its use in other settings, see [Car09, Ede13, CM21, DW22].
Neighborhood Notation: Throughout the paper, we use both open and closed neighborhoods and adopt the following notation. When $\mathcal{X}$ is a subset of a metric space $\mathcal{Z}$, through $\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})$ we will denote the open $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$ while $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathcal{Z})$ will denote the closed $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$.
@@ 2.1.1 Filtrations
As mentioned in the previous section, one needs a filtration (a nested family of topological spaces or abstract simplicial complexes) $\Delta_{\bullet}$ in order to employ the PH machinery. One of the most natural examples of such filtrations arises by considering $r$-neighborhoods, i.e., for $\mathcal{X}$ a subspace of a metric space $\mathcal{Z}$, then the family $\{\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})\}_{r\geq 0}$ defines a filtration.  While there are several ways to induce such families, from a computational perspective, a tractable method is to induce a family of simplicial complexes from a given dataset. In this paper, we are only interested in the most common ones, Vietoris–Rips complexes and Čech complexes. Note that most of our results for Čech complexes directly generalize to the alpha-complexes naturally (2.9).
@@
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a compact metric space. For $r>0$, its Vietoris–Rips complex is the abstract simplicial complex $\mathcal{V}_{r}(\mathcal{X})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{V}_{r}(\mathcal{X})$ if and only if $\mathbf{d}_{\mathcal{X}}(x_{i_{m}},x_{i_{n}})<r$ for any $0\leq m,n\leq k$.
In the sequel, given a metric space $Z$, a point $z\in Z$ and $r>0$, by $B_{r}(z)$ we will denote the open ball of radius $r$ around $z$.
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. For $r>0$, Čech complex is the abstract simplicial complex $\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ if and only if $\bigcap_{m=0}^{k}B_{r}(x_{i_{m}})\neq\emptyset$ in $\mathbf{Y}$.
Through the geometric realization functor, the nested families of simplicial complexes provided by the Vietoris-Rips and Čech complexes induce filtrations.
@@
A filtration of a topological space $Z$ is a collection $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ such that for each $0<r\leq s$, $\Delta_{r}$ is a subset of $Z$ and $\iota_{r,s}:\Delta_{r}\hookrightarrow\Delta_{s}$ is the inclusion map.
When there is no risk of confusion we will simply say that $\Delta_{\bullet}$ is a filtration without mentioning the topological space $Z$, with the understanding that $Z$ can be recovered as the colimit of $\Delta_{\bullet}$.
Example 2.4.
One example of filtrations arising in geometric scenarios is the following. Let $\mathcal{X}\subset Z$ be a nonempty compact subset of a metric space $(Z,\mathbf{d}_{Z})$. Then, one considers the filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset Z)$ given for each $r>0$ by
$\mathcal{N}_{r}(\mathcal{X}\subset Z):=\bigcup_{x\in\mathcal{X}}B_{r}(x),$
the open $r$-neighborhood of $X$ in $Z$. Of particular relevance to this paper will be the case when $Z=\mathbf{Y}$, a Banach space.
In the remainder of the paper, we use the notation $\mathcal{V}_{\bullet}(\mathcal{X})$ and $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ to denote the filtrations induced by the (geometric realizations of the) Vietoris-Rips and Čech complexes of $\mathcal{X}$, respectively. Similarly, we will denote by $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ the filtration induced by considering subsequent neighborhoods of the compact subset $\mathcal{X}$ inside of the Banach space $\mathbf{Y}$ and will refer to any such filtration arising in that manner as a neighborhood filtration.
Given that the definitions of the Vietoris-Rips and Čech simplicial complexes are similar, one can expect some relationships between these complexes. By direct computation ([EH10]), it is straightforward to see that for any $\mathcal{X}\subset\mathbf{Y}$,
$\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathcal{V}_{2r}(\mathcal{X})\subset\mathcal{C}_{2r}(\mathcal{X}\subset\mathbf{Y}).$
@@
Lemma 2.5 (Nerve Theorem [Ale28, Bor48]).
@@
There is a “persistent”, or “functorial”, version of this result.
Theorem 2.6 (Persistent Nerve Theorem [LMO23, Proposition 4.1]).
@@
[…]
Theorem 2.6 implies that the persistent homology of $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ is isomorphic to that of $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, a fact that we will repeatedly use in the sequel. Since any compact metric space $\mathcal{X}$ can be isometrically embedded into $L^{\infty}(\mathcal{X})$ (endowed with the sup norm between functions) via its Kuratowski embedding (2.23), one obtains an analogous result providing a connection between the Vietoris–Rips filtration and the nested neighborhoods $\{\mathcal{N}_{s}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subseteq\mathcal{N}_{t}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\}_{0<s<t}$ of $\mathcal{X}\subset L^{\infty}(\mathcal{X})$.
Corollary 2.7 ([LMO23, Theorem 5]).
@@
[…]
@@
The result above remains valid if $L^{\infty}(\mathcal{X})$ is substituted by any other injective metric space admitting an isometric embedding of $\mathcal{X}$; see Section 2.2 for the definition and Section 5.2, where we in particular utilize the tight span $\operatorname{E}(\mathcal{X})$ as one such choice.
@@
Note that while we only discuss VR and Čech filtered complexes in our paper, our results on Čech lifespans naturally apply to lifespans of homology classes induced by alpha complexes as $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})\simeq\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ where $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})$ represents the alpha complex induced by $\mathcal{X}$ with distance threshold $r\geq 0$; see [DW22].
@@ 2.1.2 Persistent Homology
Here, we recall basic notions pertaining to persistent homology that are necessary for our setting. For further details, see [LMO23, Section 2.1].
@@
For a given persistence module $V_{\bullet}$, if there is a multiset of intervals $(I_{\lambda})_{\lambda\in\Lambda}$ such that $V_{\bullet}$ is isomorphic to $\bigoplus_{\lambda\in\Lambda}\mathbb{F}_{\bullet}[I_{\lambda}]$, then that multiset is referred to as a (persistence) barcode associated to the persistence module $V_{\bullet}$. Persistence Modules admitting such a multiset of intervals are said to be interval decomposable. The persistence diagram of $V_{\bullet}$ is then given as the multiset of points $(b_{\lambda},d_{\lambda})\in\mathbb{R}^{2}$, where $b_{\lambda}$, is the left endpoint of $I_{\lambda}$ and $d_{\lambda}$ is its right endpoint. [^3] Not every persistence module is interval decomposable; see [CB15].
@@
For any $k\geq 0$, applying the $k$-dimensional reduced homology functor (with coefficients in a field $\mathbb{F}$) to a filtration $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ produces the persistence (homology) module $\operatorname{\tilde{H}}_{k}(\Delta_{\bullet},\mathbb{F})=(\operatorname{\tilde{H}}_{k}(\Delta_{r};\mathbb{F}),\Phi^{k}_{r,s})_{0<r\leq s}$ where the morphisms $\Phi^{k}_{r,s}$ are those induced by $\iota_{r,s}$. [^4] Note that we are using reduced homology in our definition in order to dispense with the usual infinite length bar at the level of degree zero persistent homology.
In what follows we will often drop the field $\mathbb{F}$ from the notation. Under suitable assumptions, the persistence modules obtained from filtrations, as described above, are interval decomposable. In particular, the persistence modules obtained from neighborhood filtrations of compact subsets of a Banach space are interval decomposable (so that they admit barcodes).
Theorem 2.14 ([LMO23, Theorem 1]).
Assume $\mathcal{X}$ is a compact subset of a Banach space $\mathbf{Y}$. Then there is a (unique) persistence barcode associated to the persistence module $\operatorname{\tilde{H}}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. In particular, $\operatorname{\tilde{H}}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ admits (unique) persistence barcode. [^5] In [LMO23, Theorem 1] the authors only contemplate the case of the $\mathbf{Y}$ being equal to $L^{\infty}(\mathcal{X})$ for some compact metric space $\mathcal{X}$. However, the proof of their result directly applies to the setting in the statement.
We will henceforth use $\mathrm{PD}_{k}(\mathcal{X})$ and $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ to respectively denote the persistence diagrams of the Vietoris-Rips and Čech filtrations of $\mathcal{X}$.
@@ 2.1.3 Stability Theorems
Persistence diagrams are an effective methodology for encoding topological properties of the space $\mathcal{X}$ and its neighborhoods in $\mathbf{Y}$. Persistence diagrams are stable, as expressed by the following stability theorems. Informally, the stability theorem states that if the shape and the size of two spaces are similar, then their persistence diagrams are close to each other. For a formal statement, let $\mathbf{d}_{b}(.,.)$ denote the bottleneck distance between persistence diagrams [EH10]. Let $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}$ be the Hausdorff distance between two subsets of the same Banach space $\mathbf{Y}$, and let $\mathbf{d}_{\mathrm{GH}}$ be the Gromov–Hausdorff distance between two metric spaces [BBI01].
Lemma 2.15 (Stability Theorem—VR; [${}^{+}$09, Theorem 3.1] and [CDSO14, Theorem 5.2]).
@@
Note that Lemma 2.16 implies Lemma 2.15. Indeed, this was implicity used in the proof of [${}^{+}$09, Theorem 3.1]. Notice that the coefficient $2$ does not appear in the second stability theorem.

### v2

@@ 2 Background
In this section, we provide an overview of the concepts from applied algebraic topology and metric geometry that form the foundation for the remainder of the paper. We give a summary of our notations in Table 1 in the Appendix.
@@ 2.1 Persistent Homology
Persistent Homology (PH) is a methodology rooted in Applied Algebraic Topology that captures various structural characteristics of a given topological or metric space. Its development can be traced to the pioneering work of Frosini [Fro90] and Robins [Rob99], with its algorithmic framework later established by Edelsbrunner, Letscher, and Zomorodian [ELZ02]. Earlier manifestations of persistent homology were retrospectively identified in the works of Morse [Mor30], Deheuvels [Deh55], and Barannikov [Bar94].
In the past two decades, PH has been employed as a powerful mathematical machinery for discovering patterns in data in applications within Machine Learning and Data Science. This advancement has been made possible by the development of efficient algorithms capable of computing PH in polynomial time. Specifically, the total computational effort is a polynomial function of parameters related to the size of the input simplicial filtration and the maximum homology degree to be computed; see Edelsbrunner, Letscher and Zomorodian [ELZ02], Harker, Mischaikow, Mrozek and Nanda [HMMN14, MN13], and Bauer [Bau21].
For more details on PH and its use in various settings, see Carlsson [Car09], Edelsbrunner and Harer [EH10], Chazal, de Silva, Glisse and Oudot  [CDSGO16], Ghrist [Ghr18], Rabadán and Blumberg [RB19], Carlsson and Vejdemo-Johansson [CVJ21], Joharinad and Jost [JJ23], and Polterovich, Rosen, Samvelyan and Zhang [PRSZ20].
Neighborhood Notation. Throughout the paper, we use both open and closed neighborhoods and adopt the following notation. When $\mathcal{X}$ is a subset of a metric space $\mathcal{Z}$, through
$\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z}):=\bigcup_{x\in\mathcal{X}}B_{r}(x),$
we will denote the open $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$ while $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathcal{Z})$ will denote the similarly defined closed $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$.
@@ 2.1.1 Filtrations
As noted in the previous section, utilizing the PH machinery requires a filtration—a nested family of topological spaces or abstract simplicial complexes—denoted by $\Delta_{\bullet}$. One of the most natural examples arises by considering nested neighborhoods of a subspace of a metric space, i.e., for $\mathcal{X}$ a subset of a metric space $\mathcal{Z}$, the family $\{\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})\}_{r\geq 0}$ defines a filtration. Simplicial constructions are preferred in practical applications and the most common ones are Vietoris–Rips and Čech complexes and the respective filtrations they induce. While our study focuses primarily on these two types of filtrations, most of our results concerning Čech complexes naturally extend to alpha complexes; see 2.9.
@@
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a compact metric space. For $r>0$, its Vietoris–Rips complex at scale $r$ is the abstract simplicial complex $\mathcal{V}_{r}(\mathcal{X})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{V}_{r}(\mathcal{X})$ if and only if $\mathbf{d}_{\mathcal{X}}(x_{i_{m}},x_{i_{n}})<r$ for any $0\leq m,n\leq k$.
In the sequel, given a metric space $\mathcal{Z}$, a point $z\in\mathcal{Z}$ and $r>0$, by $B_{r}(z)$ we will denote the open ball of radius $r$ around $z$.
@@
Let $\mathcal{X}$ be a compact subset of a metric space $\mathcal{Z}$. For $r>0$, the Čech complex at scale $r$ is the abstract simplicial complex $\mathcal{C}_{r}(\mathcal{X}\subset\mathcal{Z})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{C}_{r}(\mathcal{X}\subset\mathcal{Z})$ if and only if $\bigcap_{m=0}^{k}B_{r}(x_{i_{m}})\neq\emptyset$ in $\mathcal{Z}$.
In most scenarios $\mathcal{Z}=\mathbf{Y}$, a Banach space.
Through the geometric realization functor, the nested families of simplicial complexes provided by the Vietoris–Rips and Čech complexes induce filtrations.
@@
A filtration of a topological space is a collection $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ such that for each $0<r\leq s$, $\Delta_{r}$ is a subset of the topological space and $\iota_{r,s}:\Delta_{r}\hookrightarrow\Delta_{s}$ is the inclusion map.
When there is no risk of confusion we will simply say that $\Delta_{\bullet}$ is a filtration without mentioning the ambient topological space, with the understanding that it can be recovered as the colimit of $\Delta_{\bullet}$. In the remainder of the paper, we use the notation $\mathcal{V}_{\bullet}(\mathcal{X})$ and $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ to denote the filtrations induced by the (geometric realizations of the) Vietoris–Rips and Čech complexes of $\mathcal{X}$, respectively.
Example 2.4 (Neighborhod Filtrations).
Another example of filtrations arising in geometric scenarios is the following. Let $\mathcal{X}\subset\mathcal{Z}$ be a nonempty compact subset of a metric space $(\mathcal{Z},\mathbf{d}_{\mathcal{Z}})$. Then, one considers the filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathcal{Z})$ given, for each $r>0$, by the open $r$-neighborhood $\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})$ of $\mathcal{X}$ in $\mathcal{Z}$. We will refer to any filtration arising in that manner as a neighborhood filtration. Of particular relevance to this paper will be the case when $\mathcal{Z}=\mathbf{Y}$, a Banach space.
Given the similarities in the definitions of Vietoris–Rips and Čech simplicial complexes, it is natural to expect certain relationships between these two types of complexes. By direct computation, it is straightforward to see that for any compact $\mathcal{X}\subset\mathbf{Y}$,
$\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})\subseteq\mathcal{V}_{2r}(\mathcal{X})\subseteq\mathcal{C}_{2r}(\mathcal{X}\subset\mathbf{Y}).$
@@
Lemma 2.5 (Nerve Theorem; Alexandrov [Ale28] and Borsuk [Bor48]).
@@
There are “persistent”, or “functorial”, versions of this result; see Bauer, Kerber, Roll and Rolle [BKRR23] for an overview of different variants of the functorial nerve lemma. We will use the version below (see the discussion in [LMO24, Remark 4.4]) to relate the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ and the Neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$.
Theorem 2.6 (Persistent Nerve Theorem [LMO24, Proposition 4.5]).
@@
${\mathcal{C}_{s}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{C}_{t}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})}$${\mathcal{N}_{t}(\mathcal{X}\subset\mathbf{Y})}$$\scriptstyle{\varphi_{s}}$$\scriptstyle{\varphi_{t}}$
Note that Theorem 2.6 implies that the persistent homology of $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ is isomorphic to that of $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, a fact that we will repeatedly use in the sequel.
Since any compact metric space $\mathcal{X}$ can be regarded as a subset of $L^{\infty}(\mathcal{X})$ (via its Kuratowski embedding, see 2.24), one obtains an analogous result providing a connection between the Vietoris–Rips filtration and the filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ consisting of nested neighborhoods
$\big{\{}\mathcal{N}_{s}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subseteq\mathcal{N}_{t}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\big{\}}_{0<s\leq t}$
of $\mathcal{X}\subset L^{\infty}(\mathcal{X})$.
Corollary 2.7 ([LMO24, Theorem 4.1]).
@@
${\mathcal{V}_{2s}(\mathcal{X})}$${\mathcal{V}_{2t}(\mathcal{X})}$${\mathcal{N}_{s}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))}$${\mathcal{N}_{t}(\mathcal{X}\subset L^{\infty}(\mathcal{X})}$$\scriptstyle{\varphi_{s}}$$\scriptstyle{\varphi_{t}}$
@@
The result above remains valid if $L^{\infty}(\mathcal{X})$ is substituted by any other injective metric space admitting an isometric embedding of $\mathcal{X}$; see Section 2.2 for the definition and Section 5.2, where we in particular utilize the tight span $\operatorname{E}(\mathcal{X})$ as one such choice. Also, the proof of 2.7 yields that $\mathcal{C}_{\bullet}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ and $\mathcal{V}_{2\bullet}(\mathcal{X})$ are naturally homotopy equivalent; see [LMO24, Section 4].
@@
Note that while we only discuss VR and Čech filtered complexes in our paper, our results on Čech lifespans naturally apply to lifespans of homology classes induced by alpha complexes as $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})\simeq\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ where $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})$ represents the alpha complex induced by $\mathcal{X}$ with distance threshold $r\geq 0$; see Edelsbrunner and Harer [EH10, III.4].
@@ 2.1.2 Persistent Homology
Here, we recall basic notions pertaining to persistent homology that are necessary for our setting. We will follow the presentation from [LMO24, Section 2.1].
@@
For a given persistence module $V_{\bullet}$, if there is a multiset of intervals $(I_{\lambda})_{\lambda\in\Lambda}$ such that $V_{\bullet}$ is isomorphic to $\bigoplus_{\lambda\in\Lambda}\mathbb{F}_{\bullet}[I_{\lambda}]$, then that multiset is referred to as a (persistence) barcode associated to the persistence module $V_{\bullet}$. Persistence Modules admitting such a multiset of intervals are said to be interval decomposable. The persistence diagram of $V_{\bullet}$ is then given as the multiset of points $(b_{\lambda},d_{\lambda})\in\mathbb{R}^{2}$, where $b_{\lambda}$, is the left endpoint of $I_{\lambda}$ and $d_{\lambda}$ is its right endpoint. [^4] Not every persistence module is interval decomposable; see Crawley-Boevey [CB15] for more details.
@@
For any $k\geq 0$, applying the $k$-dimensional reduced homology functor (with coefficients in a field $\mathbb{F}$) to a filtration $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ produces the persistence (homology) module $\widetilde{\mathrm{H}}_{k}(\Delta_{\bullet};\mathbb{F})=(\widetilde{\mathrm{H}}_{k}(\Delta_{r};\mathbb{F}),\Phi^{k}_{r,s})_{0<r\leq s}$ where the morphisms $\Phi^{k}_{r,s}$ are those induced by $\iota_{r,s}$. [^5] Note that we are using reduced homology in our definition in order to dispense with the usual infinite length bar at the level of degree zero persistent homology.
In what follows we will drop the field $\mathbb{F}$ from the notation since all of our results hold for an arbitrary choice of $\mathbb{F}$.
Under suitable assumptions, the persistence modules obtained from filtrations, as described above, are interval decomposable. In particular, the persistence modules obtained from neighborhood filtrations of compact subsets of a Banach space are interval decomposable (so that they admit barcodes).
Theorem 2.14 ([LMO24, Theorem 1]).
Assume $\mathcal{X}$ is a compact subset of a Banach space $\mathbf{Y}$. Then there is a (unique) persistence barcode associated to the persistence module $\widetilde{\mathrm{H}}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. In particular, $\widetilde{\mathrm{H}}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ admits (unique) persistence barcode. [^6] In [LMO24, Theorem 1] the authors only contemplate the case of the $\mathbf{Y}$ being equal to $L^{\infty}(\mathcal{X})$ for some compact metric space $\mathcal{X}$. However, the proof of their result directly applies to the setting in the statement.
We will henceforth use $\mathrm{PD}_{k}(\mathcal{X})$ and $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ to respectively denote the persistence diagrams of the Vietoris–Rips and Čech filtrations of $\mathcal{X}$.
@@ 2.1.3 Stability Theorems
Persistence diagrams are an effective methodology for encoding topological properties of the space $\mathcal{X}$ and its neighborhoods in $\mathbf{Y}$. Persistence diagrams are stable, as expressed by the following stability theorems. Informally, the stability theorem states that if the shape and the size of two spaces are similar, then their persistence diagrams are close to each other. For a formal statement, let $\mathbf{d}_{b}(\cdot,\cdot)$ denote the bottleneck distance between persistence diagrams; see Edelsbrunner and Harer [EH10]. Let $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}$ be the Hausdorff distance between two subsets of the same Banach space $\mathbf{Y}$, and let $\mathbf{d}_{\mathrm{GH}}$ be the Gromov–Hausdorff distance between two metric spaces; see Burago, Burago and Ivanov [BBI01, Chapter 7].
Lemma 2.15 (Stability Theorem—VR; [CCSG+09, Theorem 3.1] and [CDSO14, Theorem 5.2]).
@@
Note that Lemma 2.16 implies Lemma 2.15. Indeed, this was implicitly used in the proof of [CCSG+09, Theorem 3.1]. Notice that the coefficient $2$ does not appear in the second stability theorem.